# Cofinality via Weighted Colimits

### v1

@@ 1 Introduction
    If $\mathscr{C}$ is an $(n,1)$-category, namely, all mapping spaces are (n1)(n-1)-truncated, then every $n$-connected space is $\mathscr{C}$-acyclic. This provides a natural interpolation between Theorem˜1.1 and Theorem˜1.3. Though most probably folklore, we are not aware of this being documented in the literature.
@@
Given functors $X\colon I\rightarrow\mathscr{C}$ and $W\colon I^{\mathrm{op}}\rightarrow\mathcal{S}$, the colimit of $X$ weighted by $W$ is the colimit of the composite functor
@@ Acknowledgments.
We are grateful to Ishai Dan-Cohen for proposing Corollary˜B as an application of a conjectural refinement of Quillen’s theorem A, and to the other participants of the Seminarak where this refinement was essentially proved, inspiring the main ideas of this note. Special thanks to Shay Ben-Moshe for useful discussions about the subject matter and to Dorin Boger for useful comments on an earlier draft. We also thank Niko Naumann for renewing our interest in Theorem˜A, encouraging us to write this note.
@@ 3 Main results
where $f^{*}$ is the restriction along $f$, and $f_{!}$ is the left Kan extension along $f$. Writing $q\colon I\rightarrow\mathrm{pt}$, we can describe $\alpha$ as the counit of the adjunction $f_{!}\dashv f^{*}$ whiskered by $q$:

### v2

@@ 1 Introduction
    If $\mathscr{C}$ is an $(n,1)$-category, namely, all mapping spaces are (n1)(n-1)-truncated, then every $n$-connected space is $\mathscr{C}$-acyclic. This provides a natural interpolation between Theorem˜1.1 and Theorem˜1.3. After an initial version of this work was completed, we became aware that this result was proved in [other_quillen_paper].
@@
Given functors $X\colon I\rightarrow\mathscr{C}$ and $W\colon I^{\mathrm{op}}\rightarrow\mathcal{S}$, the colimit of $X$ weighted by $W$, denoted $\underrightarrow{\operatorname{lim}}\,_{I}^{W}X$, is the colimit of the composite functor
@@
Let $\mathscr{C}$ be an $\infty$-category and $f\colon J\rightarrow I$ a functor. For every diagram $X\colon I\rightarrow\mathscr{C}$ and weight $W\colon J^{\mathrm{op}}\rightarrow\mathcal{S}$, we have an isomorphism
@@ Acknowledgments.
We are grateful to Ishai Dan-Cohen for proposing Corollary˜B as an application of a conjectural refinement of Quillen’s theorem A and for many useful suggestions on a previous draft. We thank the other participants of the Seminarak where this refinement was essentially proved, inspiring the main ideas of this note. Special thanks to Shay Ben-Moshe for useful discussions about the subject matter and to Dorin Boger for useful comments on an earlier draft. We also thank Niko Naumann for renewing our interest in Theorem˜A, encouraging us to write this note. The first author was supported by the DFG through SFB 1085 “Higher Invariants”.
@@ 3 Main results
where $f^{*}$ is the restriction along $f$, and its left adjoint $f_{!}$ is the left Kan extension along $f$. Writing $q\colon I\rightarrow\mathrm{pt}$, we can describe $\alpha$ as the counit of the adjunction $f_{!}\dashv f^{*}$ whiskered by $q$: